From Journal Entry to Financial Statement Bevs Dry Cleaners is owned and operated by Beverly Zahn. A building and equipment are currently being rented. The actual work of dry cleaning is done by another company and they charge a fee to Bev. The beginning balances of the businesss general ledger at January 1, 2019 are as follows: Cash debit balance $39,000 Accounts Receivable debit balance $80,000 Supplies debit balance $11,000 Land debit balance $50,000 Accounts Payable credit balance $31,500 The following transactions happened in the month of January. Please make journal entries, post to general ledger, prepare trial balance and from the trial balance prepare Balance Sheet and Income Statement for the month of January 2019 1/1 Bev Zahn invested cash of $21,000 in the business 1/2 Bev purchased land adjacent to the dry cleaner to use as a future parking lot paying cash of $35,000 1/3 the business paid rent for the month $4,000 1/5 the business charged customers for dry cleaning ON ACCOUNT, $72,000 1/10 the business paid their creditors who put the business ON ACCOUNT (the suppliers whom they owe money on) $20,0001/3 the business paid rent for the month $4,000 1/5 the business charged customers for dry cleaning ON ACCOUNT, $72,000 1/10 the business paid their creditors who put the business ON ACCOUNT (the suppliers whom they owe money on) $20,000 1/12 the business purchased supplies on account for $8,000 1/15 Received cash from customers (who are ON ACCOUNT) $35,000 1/16 Customers paid cash of $8,000 for dry cleaning jobs the Dry Clean has done for them 1/20 The business received bill from the person who is doing the actual work for the dry cleaner. The bill is not due until 2/20/2019. The total invoice amount if $29,450 1/25 paid utilities expense $1,800 1/30 The cost of supplies left is $11,800 so the business used $7200 of supplies during the month
Answer #1
Accounting Equation
Assets = Liability + Capital
Cash A/c rec Supplies land bank = A/c payab O/s Exp +
01-11-2016 39000 80000 11000 50000 = 31500 + 148500
= +
Transcation = +
A 21000 = + 21000
39000 80000 11000 50000 21000 = 31500 0 + 169500
B -35000 35000 = +
4000 80000 11000 85000 21000 = 31500 0 + 169500
C -4000 = + -4000
0 80000 11000 85000 21000 = 31500 0 + 165500
D 72000 = + 72000
0 152000 11000 85000 21000 = 31500 0 + 237500
E 0 -20000 = -20000 +
0 152000 11000 85000 1000 = 11500 0 + 237500
F 8000 = 8000 +
0 152000 19000 85000 1000 = 19500 0 + 237500
G 38000 = + 38000
38000 152000 19000 85000 1000 = 19500 0 + 275500
H 77000 -77000 = +
115000 75000 19000 85000 1000 = 19500 0 + 275500
I = 29450 + -29450
115000 75000 19000 85000 1000 = 19500 29450 + 246050
J -29200 = + -29200
85800 75000 19000 85000 1000 = 19500 29450 + 216850
K -7200 = + -7200
85800 75000 11800 85000 1000 = 19500 29450 + 209650
L -5000 = + -5000
80800 75000 11800 85000 1000 = 19500 29450 + 204650
Income Statement for Nov 16
Revenue 110000
- COGS(op+purchase-supplies) =11000+8000-11800 -7200
-Direct exp -29450
-wages -24000
Gross profit 49350
Indirect exp
- transport exp -2100
-utilities exp -1800
-rent paid -4000
-mis exp -1300
Net profit 40150
Statement of owner's Equity
Opening Capital 148500
+Addition 21000
-Withdrawal -5000
+net profit 40150
204650
Statement of Cash flow
Cash from operating activities
Net Income 40150
-Increase in CA & Decrease in CL
supplies -800
A/c payable -12000
+ Increase CL & Decrease in CA
A/c rec 5000
O/s exp 29450
Cash from operating activities 61800
Cash from investing Activites -35000
land purchased
Cash from Financing activities
Addition 21000
Withdrawal -5000 16000
Net Change in Cash 42800
Add Opening Cash 39000
Cash & Bank Balance 81800